(y+z)dx+(z+x)dy+(x+y)dz=0 find the solution



1
Expert's answer
2021-03-17T09:10:08-0400

Solve:

"(y+z)dx+(z+x)dy+(x+y)dz=0."

Solution:

Here, "P=y+z,Q=z+x,R=x+y."

"\because \sum P(\frac{\partial{Q} }{\partial{z}}-\frac{\partial{R}}{\partial{y}})=0."

condition of integrability satisfed.

Rearranging the terms of the given equation,

"(ydx+xdy)+(zdx+xdz)+(zdy+ydz)=0"

"\Rarr d(xy)+d(zx)+d(zy)=0."

on integration

Answer:

"xy+yz+zx=C."